If someone could help me factor the problem (I don't know how to write it on a computer so I am just going to write it out long hand. Bear with me here)
A squared minus B squared minus 6B minus 9
I have tried everything and can't seem to find the answer, which is (A-B-3) (A+B+3)
It has to be something totally stupid, but I cant figure it out for the life of me.
If someone could step by step it for me I would appreciate it.

Answer by khwang(438)

At first you should know hwo to type
A squared minus B squared minus 6B minus 9 as
A^2 - B^2 -6B - 9
You only need to use the basic formulas :
A^2 - B^2 = (A-B)(A+B) ...(1) and
(A+B)^2 = A^2 + 2 AB + B^2...(2)
Now A^2 - B^2 -6B - 9 = A^2 - (B^2 +6B + 9)
= A^2 - (B+3)^2 [Use formula (2), why?]
= [(A+(B+3)][A-(B+3)] [Use formula (2)]
= (A+B+3)(A-B-3)
That is the whle story.
Something if you suppose it is hard then it becomes more
difficult.However,if youthink it is easy, it becomes facile.
Math is not hard at all if you have basic confidence aboutyourself.

Normally, don't anwser such simple questions as you asked.
Here,I just don't want you see other long troublesome answers and
I don't want you give up.
Kenny